# Astrophysical Consequences of Black Holes in Asymptotically Safe Quantum Gravity: Shadows and Strong Gravitational Lensing

### v2

@@ Abstract
We investigate the astrophysical consequences of black holes in asymptotically safe quantum gravity, characterized by the parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$, in addition to the black hole mass $M$. To evaluate the physical validity of the fundamental asymptotically safe theory of quantum gravity black hole solutions, we analyze their gravitational lensing properties in the strong field regime. Specifically, we examine the shadow cast by the asymptotically safe quantum gravity black hole and constrain its parameters using observational data from the M87* and Sgr A* supermassive black holes. Our analysis reveals that, within the $1\sigma$ confidence level, a significant portion of the parameter space for asymptotically safe quantum gravity black holes is consistent with the Event Horizon Telescope (EHT) observations of M87* and Sgr A*. This suggests these black holes are plausible candidates for describing astrophysical black holes. As an additional observational test, we perform a detailed investigation of the strong gravitational lensing properties of these black holes. We explore the fundamental strong lensing observables in detail, including the angular positions and separations of the lensed images, the relative magnifications, the radius of the outermost Einstein ring, and the relativistic time delay between images. We compare the predictions of the asymptotically safe quantum gravity black hole for each observable with those of the classical Schwarzschild solution using realistic astrophysical data. Our findings provide a pathway for testing asymptotically safe quantum gravity at galactic and extragalactic scales, offering new insights into the observational properties of black hole solutions within this framework.
@@ I Introduction
The study of black holes offers a unique opportunity to test the predictions of General Relativity (GR) under extreme conditions. In the high-curvature regions surrounding black holes [1, 2], GR predicts phenomena such as time dilation, gravitational redshift, and the warping of space, which have been observed and confirmed through astrophysical measurements. The Schwarzschild and Kerr solutions to Einstein’s field equations describe the spacetime around non-rotating and rotating black holes and provide the framework for understanding black hole properties. Observing black holes is crucial in gravitational and cosmological physics, unveiling the enigmatic entities that profoundly shape spacetime. Recent developments in observational astrophysics, such as gravitational wave detection [3] from black hole mergers and direct imaging of black hole shadows [4, 5], have provided empirical support for GR in strong gravitational fields. These observations not only affirm the theoretical predictions of GR but also open new avenues for exploring potential extensions to GR, especially as researchers seek to reconcile it with quantum mechanics. Thus, black holes serve as natural laboratories, allowing scientists to probe the limits of GR and explore the fundamental nature of gravity. GR is perturbatively non-renormalizable, except for pure gravity without interactions involving scalar fields at the one-loop level [6, 7]. Consequently, GR is generally regarded as an effective field theory, posing significant challenges in developing an entirely consistent quantum theory of gravity. Nevertheless, gravity can be accurately described within a specific range of energy and length scales by imposing a suitable cutoff at high energy scales. However, beyond these limits, particularly near black hole singularities, the classical formulation of GR becomes inadequate, necessitating the development of a new physical theory.
@@
In recent years, significant advancements have been made in the field of black hole physics. However, a deeper investigation into the fascinating and complex features of black holes, particularly the formation of relativistic images, is still required. Key phenomena, including gravitational lensing [8], black hole shadows [9], and quasinormal modes [10], play a crucial role in enhancing our understanding of general relativity, modified gravity theories, quantum gravity, and quantum mechanics. Despite these advances, observing such features poses considerable challenges due to technical limitations and the need for precise theoretical modeling. Increasing attention has been directed toward understanding the impact of quantum gravity effects near singularities. Models such as M-theory, string theory, and loop quantum gravity are among the leading frameworks explored in this regard. A comprehensive formulation of quantum gravity was presented in [11], where a scale-dependent action was proposed, and an exact renormalization group equation was derived. This work marks a significant step forward in unraveling the intricate dynamics of quantum gravity in the vicinity of singularities. Even when considering GR as the low-energy limit of a more comprehensive theory known as supergravity [12], the issue of non-renormalizability remains unresolved. However, supergravity mitigates the ultraviolet (UV) divergences by reducing the number of divergent terms through its underlying supersymmetry. The quest for a consistent quantization of gravity encountered a deadlock until the advent of a non-perturbative approach to renormalization known as asymptotic safety [13, 14]. This framework posits the existence of a fixed point in the UV limit within the renormalization group flow [15, 16], allowing the gravitational coupling constant to approach this fixed point such that physical quantities become insulated from unphysical divergences. Consequently, the resulting theory is deemed asymptotically safe, implying that unphysical divergences are likely absent at high energies [17, 18]. Asymptotically safe gravity (ASG) broadens the applicability of effective field theory approaches, enabling the removal of the UV cutoff and providing a comprehensive description across all energy scales. Since ASG has profound implications at short distances (or high-energy scales) and black holes serve as natural laboratories for exploring gravity in its most extreme conditions, it is worthwhile to investigate light propagation and the gravitational deflection angle within the ASG framework. Such studies can enhance our understanding of quantum gravity from a phenomenological standpoint and provide valuable insights from an observational perspective. Furthermore, the bending of light around black holes is instrumental in shaping their shadows. Extensive research has been conducted on black hole solutions within the ASG framework [19, 20, 21, 22].
@@
Asymptotic safety, often referred to as non-perturbative renormalizability, is a concept in quantum field theory that seeks to establish a consistent and predictive quantum theory of gravity. A central aspect of this approach is the existence of a nontrivial fixed point in the renormalization group flow of the theory, which governs the behavior of coupling constants in the ultraviolet (UV) regime and ensures that physical quantities remain free from divergences. Originally proposed by Steven Weinberg as a means to formulate a theory of quantum gravity, the notion of a nontrivial fixed point as a potential UV completion applies to other field theories, particularly those that are perturbatively nonrenormalizable. In this study, we investigate the astrophysical scenario of black holes within the framework of asymptotically safe gravity [13] alongside the quantum theory of gravity, where gravitational interactions converge toward an ultraviolet fixed point at trans-Planckian scales. Over the past three decades, this quantum field theoretical approach to gravity has emerged as a prominent contender for a consistent quantum theory of gravity, particularly in the context of the functional renormalization group (FRG) method, beginning with Reuter’s seminal paper [23]. For recent reviews on this topic, see references [24, 25, 26].
@@
The study of black hole shadows and gravitational lensing further enhances our understanding of GR and its predictions around these extreme objects. The images of the supermassive black holes M87* [27, 5] and Sgr A* [28, 29] offer critical insights for testing gravitational effects in the strong field regime. A key characteristic of a black hole’s image is its shadow, formed by light rays captured by the outer event horizon. The shape and size of the shadow depend on several factors, including the parameters of the black hole, the paths followed by the light rays, and the observer’s location. When light rays approach a black hole, photons with low orbital angular momentum are trapped by the black hole, while those with higher angular momentum escape their gravitational pull. As a result, a distant observer perceives a dark region in the sky, known as the black hole’s shadow. This shadow, a silhouette of the event horizon outlined against the bright background of accreting material, is a direct observational indicator of the black hole’s presence and characteristics. The shadow’s shape and size depend on the black hole’s mass, spin, and curvature of the surrounding spacetime, offering a valuable opportunity to test the predictions of GR in the strong-field regime. For a detailed discussion of black hole shadows and their potential for testing GR, see [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
@@
Gravitational lensing has become an invaluable tool for probing fundamental properties of gravitational fields. It plays a crucial role in the study of the large-scale structure of the Universe, the behavior of massive stellar objects, and the search for dark matter candidates. In this work, we suggest that gravitational lensing can also be employed to distinguish between different gravitational theories, particularly between asymptotically safe quantum gravity and general relativity. The deflection of light by black holes and galaxies and the phenomenon of strong gravitational lensing form a significant area of research. It can potentially reveal novel astrophysical effects, providing unique signatures for testing modified theories of gravity. Observables obtained from relativistic images in strong lensing scenarios could help distinguish between different black hole metrics through high-precision observations. Gravitational lensing, a phenomenon predicted by Einstein in 1915 [47], although he initially doubted it would ever be observed. This occurs when massive objects like black holes bend and magnify light from objects behind them due to their immense gravitational fields. In the black hole regime, gravitational lensing provides a unique way to study the properties of these enigmatic objects and their influence on spacetime [48, 49, 50, 51, 52, 51, 53, 54]. When light from a distant source, such as a star or galaxy, passes close to a black hole, it can be bent into arcs or multiple images, depending on the alignment. This effect also allows us to observe a black hole’s shadow as the intense gravity bends light around the black hole, casting a silhouette against the backdrop of surrounding radiation. This effect not only magnifies distant objects but also distorts and duplicates their images, revealing information about the black hole’s mass and spin and the geometry of spacetime near the event horizon. Recent images from the Event Horizon Telescope (EHT), capturing the shadow of the supermassive black hole in the galaxy M87, have provided empirical evidence that aligns closely with GR’s theoretical predictions, showcasing both the accuracy and limitations of the theory. These observations contribute to our understanding of the gravitational dynamics around black holes while presenting new challenges and opportunities for testing GR under conditions inaccessible in any other environment. There are three primary types of gravitational lensing, classified based on the alignment of the source, lens, and observer, as well as the resulting effects on the observed light, each offering unique insights into the nature of these extreme objects. On the other hand, weak gravitational lensing involves minor distortions in the shape of background objects when light passes through a less intense gravitational field near a black hole. Microlensing occurs when a black hole or other compact object momentarily magnifies the light of a background star without creating multiple images. Strong gravitational lensing is particularly significant in the study of black holes because it produces highly visible and distinct effects, such as multiple images, arcs, or Einstein rings, which can be directly observed and analyzed. The clarity and magnitude of the distortions caused by strong lensing make it an exceptionally powerful tool for probing the fundamental nature of black holes and the behavior of light and matter in their extreme gravitational fields. The study of strong gravitational lensing around black holes has made substantial progress since Darwin’s initial work on photon orbits [55], in the Schwarzschild black hole spacetime. Building on this foundation, Virbhadra and Ellis formalized the concept of relativistic images [49], focusing on the strong-field regime where light passing near a black hole’s photon sphere can generate multiple, highly curved images. A significant breakthrough came with Bozza’s development of the strong deflection limit (SDL) [56], which provided an accurate framework for calculating light deflection in the proximity of black holes. Further advancements were made by Frittelli, Killing, and Newman, who enhanced analytical methods for solving the lens equation [57], while Bozza extended the SDL approach. This approach has been utilized across various spacetimes, including rotating and charged black holes, such as Reissner-Nordström and Kerr black holes [58, 59]. The SDL has proven to be a versatile tool, applied to classical black hole spacetimes and to modified scenarios like braneworld black holes and those arising in modified gravity theories [60, 61, 62, 63, 64, 65, 66]. Ongoing research continues to investigate gravitational lensing in diverse black hole spacetimes, including those described by higher-curvature gravity theories [67, 68, 69, 70, 67], as well as modifications of the traditional Schwarzschild geometry [71, 72, 73, 74, 75]. Several researchers have extensively explored the physical observables associated with strong lensing effects across a variety of gravitational models and black hole solutions . These studies have analyzed various astrophysical implications by examining parameters such as the angular positions of images, separations between them, magnifications, Einstein ring characteristics, and time delays in the formation of relativistic images [76, 65, 70, 77, 78, 79, 69] The present study investigates the shadow and strong gravitational lensing characteristics of black holes in asymptotically safe quantum gravity, emphasizing unique geometric properties absent in other well-known black hole solutions. Notable examples include the Van der Waals black hole [80], Kerr black hole [81], Simpson-Visser black hole [82], acoustic charged black hole [79, 83], hairy Kerr black hole [40], Kerr-Taub-NUT-Quintessence black hole [84], Kerr-Taub-NUT spacetime [85], Bardeen black hole [86], Reissner-Nordström (RN) black hole [87], and black holes from the Horndeski theory of gravity [88], among others. Gravitational lensing effects induced by various black holes have been studied both numerically and analytically in the strong-field limit [89, 90, 91]. Kumar et al. [65] analyzed the strong gravitational lensing effects of a hairy black hole in Horndeski theory of gravity. They discussed the astrophysical consequences of various supermassive black holes due to the presence of hair in this theory and compared the results with those for the standard Schwarzschild black hole in general relativity. Islam et al. [76] studied gravitational lensing by the rotating Simpson-Visser black hole in the strong-field limit. They analyzed the astrophysical implications using examples of supermassive black holes. They demonstrated that strong gravitational lensing effects can distinguish the rotating Simpson-Visser black hole from the Kerr black hole. In a subsequent study, Islam et al. [70] investigated the strong gravitational lensing effects of the Bardeen black hole within the framework of four-dimensional Einstein-Gauss-Bonnet (EGB) gravity. They explored the astrophysical consequences by examining supermassive black holes modeled with the 4D EGB Bardeen black hole and compared the results with those for the standard Schwarzschild and Bardeen black holes. Similarly, Kumar et al. [64] explored gravitational lensing in regular electrically charged (REC) black hole spacetimes, as well as REC no-horizon spacetimes, in the strong-field regime. Their analysis focused on the astrophysical implications for supermassive black holes modeled with regular spacetimes, comparing the results with those of the Schwarzschild black hole.
@@
This paper aims to extend the analysis of black hole shadows and strong gravitational lensing by incorporating the spacetime geometry of black holes in asymptotically safe quantum gravity. We explore how this promising framework, rooted in the fundamental asymptotically safe quantum gravity theory, leads to potentially observable consequences for black hole spacetimes. In particular, we focus on key lensing observables, such as the black hole shadow, deflection angle, image magnifications, and time delays, and compare them with predictions from the classical lensing scenario around a black hole in a vacuum. By analyzing the interaction between the black hole’s gravitational and quantum fields, we aim to constrain the black hole parameters through observational signatures. These results carry significant implications for understanding the environment surrounding black holes, testing GR and alternative gravity theories, and probing the fundamental properties of asymptotically safe quantum gravity through astrophysical observations.
@@
The present paper is organized as follows: Section II provides a brief overview of the black hole solutions in asymptotically safe quantum gravity, with a focus on the non-rotating case. In Section III, we derive the expression for the black hole shadow in asymptotically safe quantum gravity and present the first set of constraints on the solution parameters based on comparisons with observational data from M87* and Sgr A*. Section IV explores the strong gravitational lensing properties of the asymptotically safe quantum gravity black hole, with a detailed investigation of the associated lensing observables. Specific predictions for these observables are made using realistic astrophysical data. Finally, the conclusions and implications of our findings are discussed in Section V.
@@
In what follows, we use natural units with G = c = 1, and adopt a signature convention of {- + + +}
II Black hole solutions in asymptotically safe quantum gravity
Asymptotically Safe Gravity is a prominent and rigorous approach to Quantum Gravity, grounded in the Quantum Field Theory (QFT) framework and inspired by the Wilsonian concept of renormalization. Proposed by Weinberg in 1976, the asymptotic safety conjecture posits that gravity can become a non-perturbatively renormalizable quantum theory if the gravitational RG flow reaches an interacting fixed point in the UV regime. Asymptotically safe gravity [24, 92] is regarded as one of the most conservative and well-founded frameworks for addressing the quantum nature of gravity. It asserts that the high-energy behavior of the gravitational RG trajectory is controlled by a UV fixed point, where all relevant running couplings stabilize at finite, non-zero values. This notion of *asymptotic safety* extends the familiar concept of *asymptotic freedom*, in which couplings tend toward zero at high energies. While asymptotic freedom ensures that couplings vanish in the UV limit, asymptotic safety generalizes the idea by allowing couplings to approach constant, finite values. Both asymptotic freedom and asymptotic safety ensure that a theory is renormalizable and remains well-defined across all energy scales, providing a consistent UV-complete framework at either a free or interacting fixed point. Asymptotic safety introduces a specific scale dependence for the gravitational coupling in the ultraviolet (UV) regime, which is expected to address the curvature singularities present in classical theories. The concept of ”quantum-improving” classical black hole spacetimes by elevating the Newtonian gravitational constant, $G_{N}$, to a scale-dependent variable has garnered significant attention [93, 94, 95, 96, 97, 98]. This approach aims to capture leading-order quantum effects, yielding several intriguing implications that merit further investigation. Recent reviews have also explored these developments in detail [99, 100].

### v3

# Shadows and Strong Gravitational Lensing Around Black Hole-like Compact Object in Quadratic Gravity

@@ Abstract
We investigate the astrophysical consequences of black holes in quadratic gravity, characterized by the parameters $S_{0}$, $S_{2}$, $m_{0}$ and $m_{2}$, in addition to the black hole mass $M$. To evaluate the physical validity of the fundamental quadratic gravity black hole solutions, we analyze their gravitational lensing properties in the strong field regime. Specifically, we examine the shadow cast by the quadratic gravity black hole and constrain its parameters using observational data from the $M87^{*}$ and $SgrA^{*}$ supermassive black holes. Our analysis reveals that, within the $1\sigma$ confidence level, a significant portion of the parameter space for quadratic gravity black holes is consistent with the Event Horizon Telescope (EHT) observations of $M87^{*}$ and $SgrA^{*}$. This suggests that these black holes are plausible candidates for describing astrophysical black holes. As an additional observational test, we perform a detailed investigation of the strong gravitational lensing properties of these black holes. We explore the fundamental strong lensing observables in detail, including the angular positions and separations of the lensed images, the relative magnifications, the radius of the outermost Einstein ring and the relativistic time delay between images. We compare the predictions of the quadratic gravity black hole for each observable with those of the classical Schwarzschild solution using realistic astrophysical data. Our findings provide a pathway for testing quadratic gravity at the galactic and extragalactic scales, offering new insights into the observational properties of black hole solutions within this framework.
@@ I Introduction
The study of black holes offers a unique opportunity to test the predictions of General Relativity (GR) under extreme conditions. In the high-curvature regions surrounding black holes [1, 2], GR predicts phenomena such as time dilation, gravitational redshift, and warping of space, which have been observed and confirmed through astrophysical measurements. The Schwarzschild and Kerr solutions to Einstein’s field equations describe the spacetime around non-rotating and rotating black holes and provide the framework for understanding black hole properties. Observing black holes is crucial in gravitational and cosmological physics, unveiling the enigmatic entities that profoundly shape spacetime. Recent developments in observational astrophysics, such as gravitational wave detection [3] from black hole mergers and direct imaging of black hole shadows [4, 5], have provided empirical support for GR in strong gravitational fields. These observations not only affirm the theoretical predictions of GR, but also open new avenues for exploring potential extensions to GR, especially as researchers seek to reconcile it with quantum mechanics. Thus, black holes serve as natural laboratories, allowing scientists to probe the limits of GR and explore the fundamental nature of gravity. GR is perturbatively non-renormalizable, except for pure gravity without interactions involving scalar fields at the one-loop level [6, 7]. Consequently, GR is generally considered an effective field theory, posing significant challenges in developing an entirely consistent quantum theory of gravity. Nevertheless, gravity can be accurately described within a specific range of energy and length scales by imposing a suitable cutoff at high-energy scales. However, beyond these limits, particularly near black hole singularities, the classical formulation of GR becomes inadequate, necessitating the development of a new physical theory.
In recent years, significant advances have been made in the field of black hole physics. However, a deeper investigation of the fascinating and complex features of black holes, particularly the formation of relativistic images, is still required. Key phenomena, including gravitational lensing [8], black hole shadows [9] and quasi normal modes [10], play a crucial role in enhancing our understanding of general relativity, modified gravity theories, quantum gravity and quantum mechanics.
@@
Despite these advances, observing such features poses considerable challenges due to technical limitations and the need for precise theoretical modeling. Increasing attention has been directed toward understanding the impact of quantum gravity effects near singularities. Models such as M-theory, string theory, and loop quantum gravity are among the leading frameworks explored in this regard. A comprehensive formulation of quantum gravity was presented in [11], where a scale-dependent action was proposed and an exact renormalization group equation was derived. This work marks a significant step forward in unraveling the intricate dynamics of quantum gravity in the vicinity of singularities. Even when considering GR as the low-energy limit of a more comprehensive theory known as supergravity [12], the issue of non-renormalization remains unresolved. However, supergravity mitigates the ultraviolet (UV) divergences by reducing the number of divergent terms through its underlying supersymmetry. The quest for a consistent quantization of gravity encountered a deadlock until the advent of a non-perturbative approach to renormalization known as asymptotic safety [13, 14, 15]. This framework posits the existence of a fixed point in the UV limit within the renormalization group flow [16, 17], allowing the gravitational coupling constant to approach this fixed point so that physical quantities become insulated from unphysical divergences. Consequently, the resulting theory is deemed asymptotically safe gravity, implying that unphysical divergences are likely absent at high energies [18, 19]. To investigate how black holes are modified in quantum theory, one approach is to examine them within the framework of quadratic gravity. Unlike the perspective of effective field theory (EFT), quadratic curvature operators can emerge as fundamental modifications to General Relativity (GR), potentially inspired by quantum gravity principles. This perspective forms the basis for the theory of quadratic gravity [20]. It is crucial to distinguish between these two approaches. The inclusion of quadratic curvature operators introduces additional massive degrees of freedom [20, 21]. Specifically, in addition to the massless spin-2 degree-of-freedom characteristic of GR, these operators typically generate a massive spin-0 degree of freedom and a massive spin-2 degree of freedom. Studies of spherically symmetric black hole solutions in quadratic gravity have provided significant insights, with detailed explorations available in various works, such as[22, 23, 24, 25].
@@
In the present paper, we investigate the astrophysical scenario of black holes-like compact objects within the framework of quadratic gravity [13] alongside the quantum theory of gravity, where gravitational interactions converge toward an ultraviolet fixed point at trans-Planckian scales. Over the past three decades, this quantum field theoretical approach to gravity has emerged as a prominent contender for a consistent quantum theory of gravity, particularly in the context of the functional renormalization group (FRG) method, beginning with Reuter’s seminal paper [26]. For recent reviews on this topic,see references [27, 28, 29].
@@
The study of black hole shadows and gravitational lensing further enhances our understanding of GR and its predictions around these extreme objects. The images of the supermassive black holes $M87^{*}$ [30, 5] and $SgrA^{*}$ [31, 32] offer critical insights for testing gravitational effects in the strong field regime. A key characteristic of a black hole’s image is its shadow, formed by light rays captured by the outer event horizon. The shape and size of the shadow depend on several factors, including the parameters of the black hole, the paths followed by the light rays, and the observer’s location. When light rays approach a black hole, photons with low orbital angular momentum are trapped by the black hole, while those with higher angular momentum escape their gravitational pull. As a result, a distant observer perceives a dark region in the sky, known as the black hole’s shadow. This shadow, a silhouette of the event horizon outlined against the bright background of accreting material, is a direct observational indicator of the black hole’s presence and characteristics. The shadow shape and size depend on the black hole’s mass, spin, and curvature of the surrounding spacetime, offering a valuable opportunity to test the predictions of GR in the strong-field regime. For a detailed discussion of black hole shadows and their potential for testing GR, see [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55].
@@
Gravitational lensing has become an invaluable tool for probing fundamental properties of gravitational fields. It plays a crucial role in the study of the large-scale structure of the Universe, the behavior of massive stellar objects and the search for dark matter candidates. In this work, we suggest that gravitational lensing can also be employed to distinguish between different gravitational theories, particularly between asymptotically safe quantum gravity and general relativity. The deflection of light by black holes and galaxies and the phenomenon of strong gravitational lensing form a significant area of research. It can potentially reveal novel astrophysical effects, providing unique signatures for testing modified theories of gravity. Observables obtained from relativistic images in strong lensing scenarios could help distinguish between different black hole metrics through high-precision observations. Gravitational lensing, a phenomenon predicted by Einstein in 1915 [56], although he initially doubted it would ever be observed. This occurs when massive objects like black holes bend and magnify light from objects behind them due to their immense gravitational fields. In the black hole regime, gravitational lensing provides a unique way to study the properties of these enigmatic objects and their influence on spacetime [57, 58, 59, 60, 61, 60, 62, 63]. When light from a distant source, such as a star or galaxy, passes close to a black hole, it can be bent into arcs or multiple images, depending on the alignment. This effect also allows us to observe a black hole’s shadow as the intense gravity bends light around the black hole, casting a silhouette against the backdrop of surrounding radiation. This effect not only magnifies distant objects but also distorts and duplicates their images, revealing information about the black hole’s mass and spin and the geometry of spacetime near the event horizon. Recent images from the Event Horizon Telescope (EHT), capturing the shadow of the supermassive black hole in the galaxy $M87$, have provided empirical evidence that aligns closely with GR’s theoretical predictions, showcasing both the accuracy and limitations of the theory. These observations contribute to our understanding of the gravitational dynamics around black holes while presenting new challenges and opportunities for testing GR under conditions inaccessible in any other environment. There are three primary types of gravitational lensing, classified based on the alignment of the source, lens and observer, as well as the resulting effects on the observed light, each offering unique insights into the nature of these extreme objects. On the other hand, weak gravitational lensing involves minor distortions in the shape of background objects when light passes through a less intense gravitational field near a black hole. Microlensing occurs when a black hole or other compact object momentarily magnifies the light of a background star without creating multiple images. Strong gravitational lensing is particularly significant in the study of black holes because it produces highly visible and distinct effects, such as multiple images, arcs, or Einstein rings, which can be directly observed and analyzed. The clarity and magnitude of the distortions caused by strong lensing make it an exceptionally powerful tool for probing the fundamental nature of black holes and the behavior of light and matter in their extreme gravitational fields. The study of strong gravitational lensing around black holes has made substantial progress since Darwin’s initial work on photon orbits [64], in the Schwarzschild black hole spacetime. Building on this foundation, Virbhadra and Ellis formalized the concept of relativistic images [58], focusing on the strong-field regime where light passing near a black hole’s photon sphere can generate multiple, highly curved images. A significant breakthrough came with Bozza’s development of the strong deflection limit (SDL) [65], which provided an accurate framework for calculating light deflection in the proximity of black holes. Further advancements were made by Frittelli, Killing and Newman, who enhanced analytical methods for solving the lens equation [66], while Bozza extended the SDL approach. This approach has been utilized across various spacetimes, including rotating and charged black holes, such as Reissner-Nordström and Kerr black holes [67, 68]. The SDL has proven to be a versatile tool, applied to classical black hole spacetimes and to modified scenarios like braneworld black holes and those arising in modified gravity theories [69, 70, 71, 72, 73, 74, 75]. Ongoing research continues to investigate gravitational lensing in diverse black hole spacetimes, including those described by higher-curvature gravity theories [76, 77, 78, 79, 76], as well as modifications of the traditional Schwarzschild geometry [80, 81, 82, 83, 84]. Several researchers have extensively explored the physical observables associated with strong lensing effects across a variety of gravitational models and black hole solutions. These studies have analyzed various astrophysical implications by examining parameters such as the angular positions of images, separations between them, magnifications, Einstein rings and time delays in the formation of relativistic images have been explored extensively in [85, 74, 79, 86, 87, 88, 78]. The current study focuses on the shadow and strong gravitational lensing features of black holes-like solution in quadratic gravity, highlighting distinctive geometric properties that set them apart from other well-known black hole solutions. Notable examples include the Kerr black hole [89], the Reissner-Nordström (RN) black hole [90], the Van der Waals black hole [91], the Simpson-Visser black hole [92], the acoustic charged black hole [88], the hairy Kerr black hole [43], the Kerr-Taub-NUT spacetime [93], the Kerr-Taub-NUT-Quintessence black hole [94], the Bardeen black hole [95], and black holes arising from Horndeski gravity [96], among others.
Gravitational lensing effects caused by these black holes have been analyzed both numerically and analytically in the strong-field regime [97, 98, 99]. Kumar et al. [74] investigated the strong gravitational lensing of a hairy black hole in Horndeski gravity, discussing its astrophysical implications for supermassive black holes and comparing the results with the Schwarzschild black hole in General Relativity. Similarly, Islam et al. [85] examined the strong gravitational lensing of the rotating Simpson-Visser black hole, demonstrating its astrophysical significance and its distinct properties compared to the Kerr black hole. Islam et al. [79] analyzed the strong gravitational lensing properties of the Bardeen black hole in the context of four-dimensional Einstein-Gauss-Bonnet (EGB) gravity. They examined the astrophysical implications by modeling supermassive black holes with the 4D EGB Bardeen black hole and compared their findings to those of standard Schwarzschild and Bardeen black holes. Similarly, Kumar et al. [73] investigated gravitational lensing in the strong-field regime for regular electrically charged (REC) black hole spacetimes, as well as REC no-horizon spacetimes. Their study highlighted the astrophysical significance of supermassive black holes described by regular spacetimes, with comparisons drawn to the Schwarzschild black hole.
@@
This paper aims to extend the analysis of black hole shadows and strong gravitational lensing by incorporating the spacetime geometry of black hole-like solutions in quadratic gravity. We explore how this promising framework, rooted in the black hole-like solutions of quadratic gravity, leads to potentially observable consequences for black hole spacetimes. In particular, we focus on key lensing observables, such as the black hole shadow, deflection angle, image magnifications and time delays, and compare them with predictions from the classical lensing scenario around a black hole in a vacuum. By analyzing the interaction between the gravitational and quantum fields of the black hole, we aim to constrain the parameters of the black hole through observational signatures. These results carry significant implications for understanding the environment surrounding black holes, testing GR and alternative gravity theories, and probing the fundamental properties of quadratic gravity through astrophysical observations.
@@
The present paper is organized as follows: Section II provides a brief overview of the black hole solutions in asymptotically safe quantum gravity, with a focus on the non-rotating case. In Section III, we derive the expression for the black hole shadow in quadratic gravity and present the first set of constraints on the solution parameters based on comparisons with observational data from $M87^{*}$ and $SgrA^{*}$. Section IV explores the strong gravitational lensing properties of the quadratic gravity black hole, with a detailed investigation of the associated lensing observables. Specific predictions for these observables are made using realistic astrophysical data. Finally, the conclusions and implications of our findings are discussed in Section V.
@@
In what follows, we use natural units with G = c = 1 and adopt a signature convention of {- + + +}
II Black hole-like solutions in quadratic gravity
Quadratic gravity plays a key role in modern studies of relativistic quantum field theories. As a natural extension of Einstein’s theory, the most precise gravity framework to date, quadratic gravity introduces quadratic terms in the Lagrangian that act as corrections to general relativity, becoming significant at extremely high energies. These higher-order curvature correctios are crucial for developing a consistent quantum gravity theory, applicable near the Big Bang or spacetime singularities within black holes. Remarkably, as shown by Stelle in the 1970s [100] and reaffirmed in the literature [101, 25], the inclusion of quadratic curvature terms renders gravity renormalizable, even when coupled with a generic quantum field theory.